hi
Have a medicine that comes as a solution: 3% Aethoxysklerol medicine in 2ml.
Wish to have a concentration of 0.6% Aethoxysklerol.
How much saline would I need to add to the 2ml of 3% to get 0.6% concentration ??
Many thanks

Answer by stanbon(75887)

Have a medicine that comes as a solution: 3% Aethoxysklerol medicine in 2ml.
Wish to have a concentration of 0.6% Aethoxysklerol.
How much saline would I need to add to the 2ml of 3% to get 0.6% concentration ??
---------
Equation:
saline + saline = saline
0.97*2 + x = 0.994(x+2)
---
Multiply thru by 1000
970*2 + 1000x = 994x + 994*2
----
6x = 24*2
x = 4*2 = 8 ml (Amt of saline to add)
======================================
Cheers,
Stan H.